# Hydrogen liquid-liquid transition from first principles and machine learning

### v1

@@ Introduction.
The nature of the LLT has been debated, with many first-principles simulations [15, 22, 21, 26, 27, 30, 35] suggesting a first-order LLT below a critical temperature $T_{c}$, based on the observation of kinks in the equation of state (EOS). Given the long correlations in both time and space expected near the transition, the outcome of first-principles molecular dynamics (MD) simulations using DFT or quantum Monte Carlo (QMC) methods has been questioned because of the short time scales and/or the small system sizes considered. This could be reflected in the large variability of the predicted $T_{c}$, which ranges from $T=4000$ K [27] to about $1250$ K for the most recent simulations [30].
In the past decade, large-scale simulations have been made possible by the introduction of machine learning interatomic potentials [42] (MLIPs). These provide results at much lower computational cost, once trained on datasets generated with ab initio methods. However, the derivation of an MLIP is a delicate step per se, possibly introducing a residual bias. A recent MLIP model, based on the NequIP [43] neural network (NN) and trained on DFT data, gave a $T_{c}$ value smaller than previous estimates [44]. This model has a pressure bias of about 5 GPa in the EOS, once compared with the corresponding ab initio predictions. Recent simulations based on another MLIP trained on the same DFT functional [34] suggested that the LLT is instead a smooth crossover, even though the accuracy of the model has been criticized [35, 45].
In this Letter, we present results for the LLT obtained with MACE [46], a framework combining an equivariant message-passing NN with high-body-order messages, which shows remarkable accuracy compared to other types of MLIPs. [47] In particular, we trained a MACE MLIP on the Perdew-Burke-Ernzerhof (PBE) functional and used it to study the nature of the LLT as a function of both system size and simulation length. Our results reveal a first-order transition between an atomic liquid and a molecular solid at low temperatures ($T<950$ K), due to the melting line proximity. At higher temperatures, our simulations indicate a crossover between molecular and atomic liquids in the thermodynamic limit.
@@
where $w_{E}$, $w_{F}$, and $w_{\sigma}$ are tunable weights. However, when employing the loss in Eq. (1) to train the model  [^1] For building the model, we used $128$ equivariant messages, a correlation order of $3$, and a cutoff radius of $r_{c}=3$, we noticed the appearance of unphysical molecular solid-like structures at high temperatures ($T\sim 1200$ K) during the dynamics. Even after the inclusion of these structures in the training set, the standard loss function still energetically favored these configurations (see the Supplemental Material (SM) for details [49]). The atomic phase, on the other hand, seemed to be described correctly.
With this in mind, we modified the loss function by including a term that penalizes a global energy bias in the prediction error on molecular configurations, thus improving the description of this phase. Specifically, we used the modified loss $\mathcal{L}^{\prime}=\mathcal{L}+\Delta\mathcal{L}$ with
@@
where $\lambda$ is a parameter controlling the relative weight of the penalty with respect to the energy term of the standard loss function, and the sum only considers molecular configurations in a given training batch. To classify configurations as either atomic or molecular, we used a static criterion that solely depends on atomic positions: A configuration is classified as molecular if the first peak of its radial distribution function $g(r)$ (estimated by fitting the $g(r)$ with a Gaussian function for $r\in[0,1.3]$Å) is larger than a threshold, here set to $1.8$.
@@
To train the final MLIP, we compiled a dataset of 21 8122181221\,81221 812 configurations, partitioned into training, test, and validation sets. We extracted 17 000similar-toabsent17000$\sim$ 17\,000∼ 17 000 configurations of $128$ hydrogen atoms each from the MD simulations of Refs. [50] and [51] (see the SM for the density-temperature distribution of this set). An additional set of $3000$ 128-atoms configurations was selected from a series of AIMD simulations at lower temperatures (800 and 900 K). We also added $500$ snapshots with 256 and 512 atoms extracted from MD runs with early iterations of the MACE model. Finally, we included solid structures (see Ref. [52]) and $\sim 100$ low-temperature $96$-atoms configurations from Ref. [53]. To ensure consistency across the dataset, we recomputed energies, forces, and pressures at the PBE level, using the Quantum Espresso package [54, 55, 56]. A projector augmented-wave (PAW) pseudopotential [57] together with a $60$ Ry plane-waves cutoff was used. A sufficiently dense $\mathbf{k}$-point grid for each system size was employed to cure finite-size effects. For instance, we used a $4\times 4\times 4$ grid for the configurations with $128$ atoms.
Figure 1: Results for MD simulations obtained with the MACE model for different temperatures: (a) $T=900$ K, (b) $T=950$ K, (c) $T=1000$ K, (d) $T=1100$ K. For each temperature, the left panel reports the EOS close to the LLT for the different system sizes. For $T=900$ K the red shaded area highlights the $r_{s}$ range for which hysteresis is observed at all values of $N$. The right panels show the average $\max_{\mathbf{k}}S(\mathbf{k})/N$ along the trajectory (top right) and the average molecular fraction $m$ (bottom right) as a function of the system size $N$ for two different values of the Wigner-Seitz radius $r_{s}$.
From this dataset, we extracted $280$ configurations for testing. From the remaining configurations, $95\%$ were used for training, and $5\%$ for validation. This last group of configurations is used during the training to assess the performance of the model. At the end of the optimization, the best model is selected as the one minimizing the loss on the validation set.
@@
Table 1 reports the accuracy of this MACE model, measured by the root mean squared error (RMSE) on the energy per atom and the forces on the training, validation, and test sets. The test-set RMSE for the virial pressure is $\sim 1$ GPa. Compared to the MLIP proposed in Ref. [34], our model has an error on energy $6-7$ times smaller and is twice as accurate on forces. The recent NequIP model of Ref. [44] shows a very similar RMSE on energy, but has a $40\%$ larger RMSE on forces.
@@
To accurately study the nature of the transition, we calculated the EOS of the system on a very dense grid of Wigner-Seitz radii $r_{s}$ in the vicinity of the LLT. The results are presented in Fig. 1 for four different temperatures in the $[900-1100]$ K range. From the EOS, we can identify three distinct regimes for the LLT: At the lowest temperature computed here, i. e., $T=900$ K (Fig. 1a), the model clearly predicts a first-order transition, signaled by the hysteresis present in the EOS for all system sizes. At $T=950$ K and $1000$ K (Figs. 1b and 1c, respectively), the EOS has a strong dependence on the system size in the transition region: the smaller systems (i. e., $N=128$ and 256) suggest again a first-order transition, while for larger $N$ the pressure plateau and hysteresis are missing. Finally, at $T\geq 1100$ K (Fig. 1d) the results indicate a smooth crossover, with a relatively small size dependence.
To characterize the LLT, the sole observation of the EOS is not sufficient, since it is known that for these transitions the density is not the primary order parameter [60]. In particular, we computed the stable molecular fraction [23, 30, 47] of the system, defined in this case as the average number of hydrogen pairs whose constituent atoms stay within a distance of $1.05$ for at least a time $\tau\sim 80$ fs.
@@
Besides the stable molecular fraction, we also computed the structure factor, $S(\mathbf{k})$, to detect long-range spatial correlations, where $\mathbf{k}=\frac{2\pi}{L}(n_{1},n_{2},n_{3})$ , with $L$ being the side of the cubic simulation box and $n_{1},n_{2},n_{3}$ integer numbers. In particular, the maximum value of $S(\mathbf{k})$ can be used as a proxy for the formation of crystalline structures, with $\max_{\mathbf{k}}S(\mathbf{k})\propto N$ in a solid system in the thermodynamic limit. On the contrary, in a liquid, the rotational invariance implies that the contribution of each $\mathbf{k}$ to the $S(\mathbf{k})$ will stay constant for $N\rightarrow\infty$. The value of the average $\max_{\mathbf{k}}S(\mathbf{k})/N$ as a function of the number of particles is reported in the top right panels of Fig. 1. In the atomic phase, our simulations show that the system is a liquid at all temperatures. In the molecular region, the behavior of $\max_{\mathbf{k}}S(\mathbf{k})/N$ is instead strongly temperature dependent. Interestingly, the first-order transition observed in the EOS at $T=900$ K is accompanied by a non-vanishing value of $\max_{\mathbf{k}}S(\mathbf{k})/N$ in the thermodynamic limit, revealing a long-range spatial order of the molecular phase at this temperature (Fig. 1a). Moreover, at $T=950$ K and 1000 K, a large value of $\max_{\mathbf{k}}S(\mathbf{k})/N$ is present for $N=128$ and $256$ on the molecular side, then vanishing at larger $N$ (Figs. 1(b-c)). This suggests that the first-order transition seen in smaller systems is due to finite-size effects and it is between a molecular crystal and an atomic liquid. A genuine molecular liquid is instead observed for all system sizes at temperature $T\geq 1100$ K (Fig. 1d).
@@
Figure 2: (a) Value of $\max_{\mathbf{k}}S(\mathbf{k})/N$ as a function of the simulation time for two systems, with $N=256$ and $N=512$ atoms, respectively, at temperature $T=950$ K and $r_{s}=1.43$. Colored lines represent running averages over a time window of $8$  ps. (b) Results for $200$ ps long simulations, $N=128$, $T=1400$ K and corresponding confidence intervals, estimated from a running average with time window $\tau_{\textrm{run}}=4$ ps. The AIMD result reported in Ref. [50] is also shown. (c) Same as (b) but with $N=512$, $T=1000$ K and $\tau_{\textrm{run}}=10$ ps.
This is further demonstrated in Fig. 2a, where the behavior of $\max_{\mathbf{k}}S(\mathbf{k})/N$ is plotted as a function of the simulation time for two sizes, namely $N=256$ and 512, at $T=950$ K and $r_{s}=1.43$. While the smaller system is in a solid-like state for the majority of the time, the system with $N=512$ occasionally crystallizes, as indicated by the two jumps of $\max_{\mathbf{k}}S(\mathbf{k})/N$ at the beginning and the end of the MD run, but it mostly remains liquid. This behavior not only confirms the size dependence already seen, but also shows that at least $10$ ps are necessary to melt the crystal for $N=512$, suggesting that $\sim 100$ ps are needed to have truly converged results near the transition.
Figure 3: Specific heat per particle vs pressure along the isotherms, for $N=2048$ obtained with the MACE model. The shaded areas indicate the uncertainty in the peak position. In the inset, we report the size scaling of $c_{v}/k_{B}N$ for four different temperatures. The results confirm a first-order transition for $T=900$ K and a smooth crossover for higher temperatures in the thermodynamic limit.
To analyze the possibility that the kinks identified in the EOS of previous AIMD results originate from the short length of these simulations, we analyzed $0.2$ ns long trajectories obtained with the MACE model for values of $r_{s}$ close to the transition at different temperatures above our $T_{c}$. The results are shown in Figs. 2b and 2c for the system of $N=128$ at $T=1400$ K and the one of $N=512$ at $T=1000$ K, respectively. Using long trajectories, we computed the pressure running average using a variable-size time window $\tau_{\textrm{run}}$, corresponding to simulations times achievable by AIMD, e. g., $\tau_{\textrm{run}}\sim 10$ ps. The variance of the running average gives an idea of the variability of the estimated equilibrium pressure when running MD simulations of length equal to $\tau_{\textrm{run}}$. From Figs. 2b and 2c, we can notice that the estimated variance grows in the vicinity of the transition, and that the presence of a pressure plateau in the EOS can be understood as an artifact due to the lack of phase-space sampling.
@@
Both the first-order transition and the crossover above the critical temperature $T_{c}$ can be further characterized by studying the behavior of the isochoric specific heat (i. e., the heat capacity per particle), $c_{v}=\frac{k_{B}}{NT^{2}}(\left\langle E^{2}\right\rangle-\left\langle E\right\rangle^{2})$, as a function of temperature, density, and system size. The results for $c_{v}$ obtained from simulations with $N=2048$ hydrogen atoms at temperatures up to $1800$ K are reported in Fig. 3. For each temperature, we estimated the location and value of the maximum. For temperatures between $900$ K and $1100$ K, we also studied the size scaling of the $c_{v}$ maximum, as shown in the inset of Fig. 3. In a first-order transition, the $c_{v}$ peak presents a divergence scaling like $\sim N$ in the thermodynamic limit [61]. As shown in the figure, linear scaling is observed at $T=900$ K, while the ratio $c_{v}/N$ tends to zero for higher temperatures. This is consistent with our EOS results locating the critical point between $T=900$ K and $T=950$ K. The position of the $c_{v}$ maximum at higher temperatures allows one to determine the Widom line in the supercritical region. The results are reported in Fig. 4. Our Widom line shows remarkable agreement with previous simulations of the LLT obtained with the PBE functional. In particular, our results well reproduce the ones reported in Ref. [21] for temperatures up to $1400$ K and those of Ref. [35] for higher temperatures, even though they do not agree on the LLT first-order character for $T\geq 900$ K.
Figure 4: Classical PBE-LLT location as computed with different methods. AIMD results by Ref. [21] (light blue dashed line), Ref. [22] (green dashed line), Ref. [35] (blue dashed line). The results for the molecular-to-atomic crossover obtained with an NN MLIPs by Ref. [34] are reported with orange and violet markers, corresponding to the maximum of the isobaric specific heat and density, respectively. The black markers and line indicate the recently proposed PBE melting line by Ref.[53], obtained with an NN MLIP and the two-phase method. The MACE model results are indicated with red markers. The filled point at $T=900$ K indicates the first-order character of the transition, while the empty points correspond to the location of the Widom line given by the $c_{v}$ maximum (see Fig. 3).

### v2

@@ Introduction.
The nature of the LLT has been debated, with many first-principles simulations [15, 22, 21, 26, 27, 30, 35] suggesting a first-order LLT below a critical temperature $T_{c}$ and pressure $p_{c}$, based on the observation of kinks in the equation of state (EOS). Given the long correlations in both time and space expected near the transition, the outcome of first-principles molecular dynamics (MD) simulations using DFT or quantum Monte Carlo (QMC) methods has been questioned because of the short time scales and/or the small system sizes considered. This could be reflected in the large variability of the predicted $T_{c}$ and $p_{c}$, which range from $T_{c}\sim 4000$ K and $p_{c}\sim 30$ GPa [27, 29], to $T_{c}\sim 1250$ K and $p_{c}\sim 150$ GPa for the most recent simulations [30].
In the past two decades, large-scale simulations have been made possible by the introduction of machine learning interatomic potentials [42] (MLIPs). These provide results at much lower computational cost, once trained on datasets generated with ab initio methods. However, the derivation of an MLIP is a delicate step per se, possibly introducing a residual bias. A recent MLIP model, based on the NequIP [43] neural network (NN) and trained on DFT data, gave a $T_{c}$ value smaller than previous estimates [44]. This model has a pressure bias of about 5 GPa in the EOS, once compared with the corresponding ab initio predictions. Recent simulations based on another MLIP trained on the same DFT functional [34] suggested that the LLT is instead a smooth crossover, even though the accuracy of the model has been criticized [35, 45].
In this work, we present results for the LLT obtained with MACE [46], a framework combining an equivariant message-passing NN with high-body-order messages, which shows remarkable accuracy compared to other types of MLIPs. [47] In particular, we trained a MACE MLIP on the Perdew-Burke-Ernzerhof (PBE) functional and used it to study the nature of the LLT as a function of both system size and simulation length. This also allows us to directly compare our results with the large body of literature that used the same functional [1, 18, 19, 21, 22, 23, 25, 27, 29, 30, 33, 34, 35, 44, 37, 45, 47]. Our results reveal a first-order transition between an atomic liquid and a molecular solid at low temperatures ($T<950$ K), due to the melting line proximity. At higher temperatures, our simulations indicate a crossover between molecular and atomic liquids in the thermodynamic limit.
@@
where $w_{E}$, $w_{F}$, and $w_{\sigma}$ are tunable weights. However, when employing the loss in Eq. (1) to train the model  [^1] For building the model, we used $128$ equivariant messages, a correlation order of $3$, and a cutoff radius of $r_{c}=3$ Å, we noticed the appearance of molecular solid-like structures at high temperatures ($T\sim 1200$ K) during the dynamics. AIMD simulations at the same conditions and system sizes show how these configurations are not dynamically stable, and, therefore, are an artifact of the model. Even after the inclusion of these structures in the training set, the standard loss function still energetically favored these configurations (see the Supplemental Material (SM) for details [49]). The atomic phase, on the other hand, seemed to be described correctly.
An analysis of the partial energy error distributions for this model, computed on subsets of the training set describing the selected phases, reveals how the loss function $\mathcal{L}$ produces a bias, i.e., a non-zero mean of $E^{\textrm{pred}}_{\mu}-E^{\textrm{ref}}_{\mu}$. Indeed, being a function of the MSEs, $\mathcal{L}$ only guarantees that the total error distribution is centered around zero.
With this in mind, we modified the loss function by including a term that penalizes a global energy bias in the prediction error on the subset of molecular configurations, thus improving the description of this phase. Specifically, we used the modified loss $\mathcal{L}^{\prime}=\mathcal{L}+\Delta\mathcal{L}$ with
@@
where $\lambda$ is a parameter controlling the relative weight of the penalty with respect to the energy term of the standard loss function, $w_{E}$ is the same hyperparameter as in Eq. (1), and the sum only considers molecular configurations in a given training batch. To classify configurations as either atomic or molecular, we used a static criterion that solely depends on atomic positions: A configuration is classified as molecular if the first peak of its radial distribution function $g(r)$ (estimated by fitting the $g(r)$ with a Gaussian function for $r\in[0,1.3]$Å) is larger than a threshold, here set to $1.8$.
@@
To train the final MLIP, we compiled a dataset of 21 81221\,812 configurations. We extracted $\sim 17\,000$ configurations of $128$ hydrogen atoms each from the MD simulations of Refs. [50] and [51] (see the SM for the density-temperature distribution of this set). An additional set of $3000$ 128-atoms configurations was selected from a series of AIMD simulations at lower temperatures (800 and 900 K). We also added $500$ snapshots with 256 and 512 atoms extracted from MD runs with early iterations of the MACE model. Finally, we included solid structures (see Ref. [52]) and $\sim 100$ low-temperature $96$-atoms configurations from Ref. [53]. To ensure consistency across the dataset, we recomputed energies, forces, and pressures at the PBE level, using the Quantum Espresso package [54, 55, 56]. A projector augmented-wave (PAW) pseudopotential [57] together with a $60$ Ry plane-waves cutoff was used. A sufficiently dense $\mathbf{k}$-point grid for each system size was employed to cure finite-size effects. For instance, we used a $4\times 4\times 4$ grid for the configurations with $128$ atoms.
Figure 1: Results for MD simulations obtained with the MACE model for different temperatures: (a) $T=900$ K, (b) $T=950$ K, (c) $T=1000$ K, (d) $T=1100$ K. For each temperature, the left panel reports the EOS close to the LLT for the different system sizes. For $T=900$ K the red shaded area highlights the $r_{s}$ range for which hysteresis is observed at all values of $N$. The right panels show the average $\max_{\mathbf{k}}S(\mathbf{k})/N$ along the trajectory (top right) and the average molecular fraction $m$ (bottom right) as a function of the system size NN for two different values of the Wigner-Seitz radius $r_{s}$.
From this dataset, we randomly extracted $280$ configurations for testing. From the remaining configurations, $95\%$ were used for training, and $5\%$ for validation. This last group of configurations is used during the training to assess the performance of the model. At the end of the optimization, the best model is selected as the one minimizing the loss on the validation set.
@@
Table 1 reports the accuracy of this MACE model, measured by the root mean squared error (RMSE) on the energy per atom and the forces on the training, validation, and test sets. The test-set RMSE for the virial pressure is $\sim 1$ GPa. Compared to the MLIP proposed in Ref. [34], our model has an error on energy 676-7 times smaller and is twice as accurate on forces. The recent NequIP model of Ref. [44] shows a very similar RMSE on energy, but has a $40\%$ larger RMSE on forces.
@@
To accurately study the nature of the transition, we calculated the EOS of the system in the vicinity of the LLT, using a dense grid of Wigner-Seitz radii $r_{s}$, where $\frac{4\pi}{3}\left(r_{s}a_{0}\right)^{3}=\frac{V}{N}$ with $V$ and $a_{0}$ being the system volume and the Bohr radius, respectively. The results are presented in Fig. 1 for four different temperatures in the [9001100][900-1100] K range. From the EOS, we can identify three distinct regimes for the LLT: At the lowest temperature computed here, i. e., $T=900$ K (Fig. 1a), the model clearly predicts a first-order transition, signaled by the hysteresis present in the EOS for all system sizes. At $T=950$ K and $1000$ K (Figs. 1b and 1c, respectively), the EOS has a strong dependence on the system size in the transition region: the smaller systems (i. e., $N=128$ and 256) suggest again a first-order transition, while for larger $N$ the pressure plateau and hysteresis are missing. Finally, at $T\geq 1100$ K (Fig. 1d) the results indicate a smooth crossover, with a relatively small size dependence.
To characterize the LLT, the sole observation of the EOS is not sufficient, since it is known that for these transitions the density is not the primary order parameter [60]. In particular, we computed the stable molecular fraction [23, 30, 47] of the system, defined in this case as the average number of hydrogen pairs whose constituent atoms stay within a distance of $1.05$ Å for at least a time $\tau\sim 80$ fs.
@@
Besides the stable molecular fraction, we also computed the structure factor to detect long-range spatial correlations, i.e.,
$S(\mathbf{k})=\left\langle\frac{1}{N}\sum_{i,j}\exp\bigl(i\mathbf{k}\cdot\left(\mathbf{R}_{i}-\mathbf{R}_{j}\right)\bigr)\right\rangle,$
where $\langle\cdots\rangle$ indicates the average over the trajectory, $\mathbf{R}_{1},\dots,\mathbf{R}_{N}$ are atom positions, and $\mathbf{k}=\frac{2\pi}{L}(n_{1},n_{2},n_{3})$, with $L$ being the side of the cubic simulation box and $n_{1},n_{2},n_{3}$ integer numbers. In particular, the maximum value of $S(\mathbf{k})$ can be used as a proxy for the formation of crystalline structures, with $\max_{\mathbf{k}}S(\mathbf{k})\propto N$ in a solid system in the thermodynamic limit. On the contrary, in a liquid, the rotational invariance implies that the contribution of each $\mathbf{k}$ to the $S(\mathbf{k})$ will stay constant for $N\rightarrow\infty$. The value of the average $\max_{\mathbf{k}}S(\mathbf{k})/N$ as a function of the number of particles is reported in the top right panels of Fig. 1. In the atomic phase, our simulations show that the system is a liquid at all temperatures. In the molecular region, the behavior of $\max_{\mathbf{k}}S(\mathbf{k})/N$ is instead strongly temperature dependent. Interestingly, the first-order transition observed in the EOS at $T=900$ K is accompanied by a non-vanishing value of $\max_{\mathbf{k}}S(\mathbf{k})/N$ in the thermodynamic limit, revealing a long-range spatial order of the molecular phase at this temperature (Fig. 1a). Moreover, at $T=950$ K and 1000 K, a large value of $\max_{\mathbf{k}}S(\mathbf{k})/N$ is present for $N=128$ and $256$ on the molecular side, then vanishing at larger $N$ (Figs. 1(b-c)). This suggests that the first-order transition seen in smaller systems is due to finite-size effects and it is between a molecular crystal and an atomic liquid. A genuine molecular liquid is instead observed for all system sizes at temperature $T\geq 1100$ K (Fig. 1d).
@@
Figure 2: (a) Value of $\max_{\mathbf{k}}S(\mathbf{k})/N$ as a function of the simulation time for two systems, with $N=256$ and $N=512$ atoms, respectively, at temperature $T=950$ K and $r_{s}=1.43$. Colored lines represent running averages over a time window of 88  ps. (b) Results for $200$ ps long simulations, $N=128$, $T=1400$ K and corresponding confidence intervals, estimated from a running average with time window $\tau_{\textrm{run}}=4$ ps. The AIMD result reported in Ref. [50] is also shown. (c) Same as (b) but with $N=512$, $T=1000$ K and $\tau_{\textrm{run}}=10$ ps.
This is further demonstrated in Fig. 2a, where the behavior of $\max_{\mathbf{k}}S(\mathbf{k})/N$ is plotted as a function of the simulation time for two sizes, namely $N=256$ and 512, at $T=950$ K and $r_{s}=1.43$. While the smaller system is in a solid-like state for the majority of the time, the system with $N=512$ occasionally crystallizes, as indicated by the two jumps of $\max_{\mathbf{k}}S(\mathbf{k})/N$ at the beginning and the end of the MD run, but mostly remains liquid. This behavior not only confirms the size dependence already seen, but also shows that at least $10$ ps are necessary to melt the crystal for $N=512$, suggesting that $\sim 100$ ps are needed to obtain fully converged results near the transition.
Figure 3: Specific heat per particle vs pressure along the isotherms, for $N=2048$ obtained with the MACE model. The shaded areas indicate the uncertainty in the peak position. In the inset, we report the size scaling of the maximum of $c_{v}/k_{B}$ for four different temperatures (in log-log scale). The results confirm a first-order transition for $T=900$ K, where $c_{v}$ scales with a behavior close to linear, and a smooth crossover for higher temperatures in the thermodynamic limit.
To analyze the possibility that the kinks identified in the EOS of previous AIMD results originate from the short length of these simulations, we analyzed $0.2$ ns long trajectories obtained with the MACE model for values of $r_{s}$ close to the transition at different temperatures above our $T_{c}$. The results are shown in Figs. 2b and 2c for the system of $N=128$ at $T=1400$ K and the one of $N=512$ at $T=1000$ K, respectively. Using long trajectories, we computed the pressure running average using a variable-size time window $\tau_{\textrm{run}}$, corresponding to simulation times achievable by AIMD, e. g., $\tau_{\textrm{run}}\sim 10$ ps. The variance of the running average gives an idea of the variability of the estimated equilibrium pressure when running MD simulations of length equal to $\tau_{\textrm{run}}$. From Figs. 2b and 2c, the estimated variance grows in the vicinity of the transition, and the presence of a pressure plateau in the EOS can be understood as an artifact due to the lack of phase-space sampling.
@@
Both the first-order transition and the crossover above the critical temperature $T_{c}$ can be further characterized by studying the behavior of the isochoric specific heat (i. e., the heat capacity per particle), $c_{v}=\frac{1}{Nk_{B}T^{2}}(\left\langle E^{2}\right\rangle-\left\langle E\right\rangle^{2})$, as a function of temperature, density, and system size. The results for $c_{v}$ obtained from simulations with $N=2048$ hydrogen atoms at temperatures up to $1800$ K are reported in Fig. 3. For each temperature, we estimated the location and value of the maximum. For temperatures between $900$ K and $1100$ K, we also studied the size scaling of the $c_{v}$ maximum, as shown in the inset of Fig. 3. In a first-order transition, the $c_{v}$ peak presents a divergence scaling like $\sim N$ in the thermodynamic limit [61]. As shown in the figure, linear scaling is observed at $T=900$ K, while $c_{v}$ tends to a constant for higher temperatures. This is consistent with our EOS results locating the critical point between $T=900$ K and $T=950$ K. The position of the $c_{v}$ maximum at higher temperatures allows one to determine the Widom line in the supercritical region. The results are reported in Fig. 4. Our Widom line shows remarkable agreement with previous simulations of the LLT obtained with the PBE functional. In particular, our results well reproduce the ones reported in Ref. [21] for temperatures up to $1400$ K and those of Ref. [35] for higher temperatures, even though they do not agree on the LLT first-order character for $T\geq 900$ K.
Figure 4: Classical PBE-LLT location as computed with different methods. AIMD results by Ref. [21] (light blue dashed line), Ref. [22] (green dashed line), Ref. [35] (blue dashed line). The results for the molecular-to-atomic crossover obtained with an NN MLIP by Ref. [34] are reported with orange and violet markers, corresponding to the maximum of the isobaric specific heat and density, respectively. The black markers and line indicate the recently proposed PBE melting line by Ref. [53], obtained with an NN MLIP and the two-phase method. The MACE model results are indicated with red markers. The filled point at $T=900$ K indicates the first-order character of the transition, while the empty points correspond to the location of the Widom line given by the $c_{v}$ maximum (see Fig. 3).